OpenStudy (anonymous):

Feel stupid can't remember how to work out this problem 4(5n-7)=32

OpenStudy (xishem):

\[4(5n-7)=32\]\[20n-28=32\]\[20n=60\]\[n=3\]

OpenStudy (anonymous):

Thank you trying to help my little sister and could not get it. Thanks I remember now.

OpenStudy (anonymous):

\[4(5n-7)=32\] \[5n-7=8\] \[5n=15\] \[m=3\]